# Novel Lower Bounds on $M/G/k$ Scheduling

### v2

@@ 8. Deriving the Test Functions – DiffeDrift
We start by deriving the affine-drift test function (Definition 7.2) in Section 8.1 as well as the constant-drift test function (Definition 7.1), which we used to characterize the mean total work of an ISQ-2 system. We then derive the modified affine-drift test function (Definition 7.5) in Section 8.2, which is a specialized test function for the AR-ISQ-2 system.
@@ 8.1. Affine-drift and constant-drift test functions
which has a unique solution given by $h_{2}(w,1/2)=w^{2}-\frac{1-e^{-2w\lambda}}{2\lambda^{2}}+\frac{w}{\lambda}.$ Ssolving this differential equation is the essence of our Diffedrift method. More complex in general $k$ case. This defines the affine-drift test function $h_{2}(w,i)$ over the three possible states $(0,0),(w,1/2)$ and $(w,1)$, as given in Definition 7.2.
@@ 8.2. Modified affine-drift test function
The affine-drift and constant-drift test functions are sufficient to determine the mean work in the ISQ-2 system. They also suffice to determine the mean relevant work in the Sep-ISQ-2 system, in the AR-ISQ-2 system an recycling stream with jobs of size $x$ arrives into the same system as the truncated stream. Therefore, $\mathbb{P}(I=0)$ is no longer given by the expression derived in the proof of Proposition 7.3 and we cannot apply the result of Proposition 7.3 in determining the changes in the drift of the AR-ISQ-2 system due to the truncated stream. Letting $I^{r}$ denote the speed distribution in $\mathbb{P}(I^{r}=0)$, which now depends on the specific recycling stream, which is difficult to characterize exactly.
@@
This is a characterization of the unused capacity in the AR-ISQ-2 system, which is unaffected by the details of recycling stream. Therefore, our plan is to modify the affine-drift test function $h_{2}$ so that we get an unused-load term matching (8.1). Additionally, we use Lemma 5.6 instead of Lemma 5.5 because it supports recyclings. We start with the following test functions:
@@
In particular, we want to ensure that (8.2) is a constant not depending on $w$. By doing so, we will create a term matching the unused-load (8.1), allowing us to effectively bound the AR-ISQ-2 system. We call this constant $C_{2}(x,\lambda_{x})$, though we do not yet know its exact value. The drift at speed $1/2$ is given by,
@@
We were able to apply the unused capacity formula (8.1) because our test function satisfied (8.2). This is key to the strength of our AR-ISQ-2 bound Theorem 7.6.
@@ 9. Bounding Mean Relevant Work in the $M/G/k$
In this section, we extend the results in Section 7 for the ISQ-2 system to the general ISQ-$k$ system. The derivation of the test functions follow the same ideas outlined in Section 8. We start by deriving our test functions using the DiffeDrift method. We first define the ISQ-$k$ constant-drift test function.
@@
[…]
@@
[…]
@@
Note that for $k=2$ this simplifies to the $h_{2}$ expression defined in Definition 7.2 and that the $v_{q}(w)$ formulas have a similar recursive structure as $u_{q}(w)$, but with an additional $y$ coefficient in the numerator of the integrand fraction. Equations 9.1 and 9.2 can be explicitly solved for any $k$, see Appendix C for the three server case. In Lemmas B.1 and B.3 of Appendix B we prove the validity of these test functions.
@@
For an arbitrary job size distribution $S$ such that $E[S^{3}]$ is finite and arrival rate $\lambda$. The expected total work in an ISQ-$k$ system is given by
@@
By Lemma B.2 and Lemma B.4 in Appendix B the drifts of the constant and affine test functions at speed 00 are given by $G\circ g_{k}(0,0)=\lambda\mathbb{E}[S+u_{1}(S)]$ and $G\circ h_{k}(0,0)=\lambda\mathbb{E}[S^{2}+v_{1}(S)]$ respectively. At all other speeds $i\geq 1/k$, the drift of the constant-drift test function is given by $G\circ g_{k}(w,i)=\lambda\mathbb{E}[S]-1$ and the drift of the affine-drift test function is given by $G\circ h_{k}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(\lambda\mathbb{E}[S]-1)$.
@@
Next, using Proposition 9.3 we give the exact mean relevant work in the Sep-ISQ-$k$ system. This constitutes the first lower bound of Theorem 4.2.
@@
For an arbitrary job size distribution $S$ and arrival rate $\lambda$. The expected relevant work in the separate-recycling-ISQ-k system is given by,
$\mathbb{E}[W_{x}^{\text{sep-ISQ}}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}+\dfrac{k\lambda(1-F_{S}(x))x^{2}}{2},$ (9.5)
where $v_{1}$ and $u_{1}$ are defined in Equations 9.1 and 9.2.
@@
The relevant work in the separate-recycling-ISQ-$k$ system is the sum of the total work in the truncated-ISQ-$k$ and the relevant work in the separate $M/G/\infty$ system.
@@
For the last part of this section, we derive the AR-ISQ-$k$ lower bound. We start with the ISQ-$k$ modified affine-test function, denoted by $h_{k,x}$.
@@
[…]
@@
where $v_{1}$ and $u_{1}$ are defined by Equations 9.1 and 9.2.
@@
For an arbitrary job size distribution $S$ and arrival rate $\lambda$. A lower bound on expected relevant work in the arbitrary-recycling-ISQ-$k$ system is given by,
$\mathbb{E}[W_{x}^{\text{arb-ISQ}}]\geq\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}\cdot\dfrac{1-\rho_{\bar{x}}}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})J_{x}}{2(1-\rho_{x})}.$ (9.8)
@@
We believe that $J_{x}=x^{2}$ for all thresholds $x$, numbers of servers $k$, and job size distributions $S$. We proved in Theorem 7.6 that $J_{x}=x^{2}$ whenever $k=2$. However, formal proof of this assertion for $k\geq 3$ remains elusive, so we leave this as an open problem. We have empirically verified that this is true for the 3-server setting with exponential job sizes.
@@
For any job size distribution $S$ and any number of servers $k\geq 2,J_{x}=x^{2}$.
10. WINE Bounds
To translate lower bounds on the mean relevant work of the $M/G/k$ system into lower bounds on its mean response time under arbitrary scheduling policies, we use the Work Integral Number Equality (WINE) introduced in Theorem 6.3 of [32]. WINE allows us to characterize the mean response time of a generic queueing system under a generic scheduling policy in terms of the mean relevant work for each relevancy-cutoff level $x$.
Proposition 10.1 (WINE Identity).
For an arbitrary scheduling policy $\pi$, in an arbitrary system,
[…]
WINE has been used to upper bound mean response time under complex scheduling policies in the $M/G/1$ [33] as well as in multiserver systems ([32] and [12]).
The first lower bound is the resource-pooled $M/G/1$ queue under the single-server-optimal SRPT policy. The optimal policy in the resource-pooled $M/G/1$ must lower bound the optimal policy in the $M/G/k$, as the resource-pooled $M/G/1$ is strictly more flexible, and could emulate the $M/G/k$ policy if desired. [29] shows that SRPT is optimal for $M/G/1$, so it must lower bound the mean response time of the $M/G/k$ system. [9] proves that this bound is tight at high load, but it is empirically loose at medium and low load (see Figure 1).
[30] characterizes the mean relevant work in the SRPT-1 system:
The second bound is the mean service time bound. To compute the mean relevant work, we imagine jobs arriving into an $M/G/\infty$ queue, whose servers run at the same speed as the $M/G/k$ servers. Each job of initial size $s$ will spend $k\min\{s,x\}$ amount of time as a relevant job, as its size shrinks linearly down to 00. The job thus contribute $\frac{k}{2}\min\{s,x\}^{2}$ amount of relevant work. Therefore,
Taking the maximum of these two bounds for each $x$ and integrating using WINE (Proposition 10.1) gives us our first improved lower bound on the mean response time of the $M/G/k$ under an arbitrary scheduling policy, which we call the WINE-2 bound.
We further improve upon WINE-2 by incorporating the exact mean relevant work of the Sep-ISQ-$k$ system, giving an improved lower bound on mean relevant work in the $M/G/k$, namely Equation 4.3. Combining these bounds with WINE results in a new lower bound which we call WINE-3. Finally, WINE-3 is augmented with a lower bound on the mean relevant work of the AR-ISQ-$k$ system, giving a final lower bound on the $M/G/k$, namely Equation 4.4, resulting in the final WINE-4 bound.
10.1. Numerical Comparison of Bounds
We illustrate the impact of each of these new lower bounds (Equations 4.1 and 4.2 of Theorem 4.1) in our WINE lower bounding framework at different thresholds $x$ in Figure 2. Specifically, we show the weighted expected relevant work $\mathbb{E}[W_{x}]/x^{2}$ for each lower bound in an $M/G/2$ with exponential job size distribution setting. We plot the weighted expected relevant work lower bounds given by (4.1) and (4.2) against the $M/G/\infty$ and resource-pooled $M/G/1/SRPT$ queues. The job size distribution is $S\sim Exp(1)$ and the arrival rate is $\lambda=0.9$.
We observe that $\mathbb{E}[W_{x}^{M/G/\infty}]$ and $\mathbb{E}[W_{x}^{sep-ISQ}]$ are almost identical for small values of $x$, but as $x$ increases, $\mathbb{E}[W_{x}^{sep-ISQ}]$ eventually overtakes $\mathbb{E}[W_{x}^{M/G/\infty}]$. Then, for large values of $x$, $\mathbb{E}[W_{x}^{arb-ISQ}]$ surpasses all of the other bounds. This occurs because, as $x$ increases, the factor $\frac{1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}}{1-\rho_{x}}$ in the second term of (4.2) approaches $1$, while the last term scales with the inverse of $1-\rho_{x}$. This contrasts with (4.1), where such scaling behavior is missing.
Figure 2. $\mathbb{E}[W_{x}]/x^{2}$ for four lower bounds on mean relevant work in the $M/G/2$ with $S\sim Exp(1)$ job size distribution and arrival rate $\lambda=0.9.$ The blue and orange curves are our novel lover bounds while the dashed red and green curves are naive lower bounds.

### v3

@@ 9. Deriving the Test Functions – DiffeDrift
We start by deriving the affine-drift test function (Definition 8.2) in Section 9.1 as well as the constant-drift test function (Definition 8.1), which we used to characterize the mean total work of an ISQ-2 system. We then derive the modified affine-drift test function (Definition 8.5) in Section 9.2, which is a specialized test function for the Rec-ISQ-2 system.
@@ 9.1. Affine-drift and constant-drift test functions
which has a unique solution given by $h_{2}(w,1/2)=w^{2}-\frac{1-e^{-2w\lambda}}{2\lambda^{2}}+\frac{w}{\lambda}.$ Solving this differential equation is the essence of our DiffeDrift method. This defines the affine-drift test function $h_{2}(w,i)$ over the three possible states $(0,0),(w,1/2)$ and $(w,1)$, as given in Definition 8.2.
@@ 9.2. Modified affine-drift test function
The affine-drift and constant-drift test functions are sufficient to determine the mean work in the ISQ-2 system. They also suffice to determine the mean relevant work in the Sep-ISQ-2 system. However, in the Rec-ISQ-2 system an recycling stream with jobs of size $x$ arrives into the same system as the truncated stream. Therefore, $\mathbb{P}(I=0)$ is no longer given by the expression derived in the proof of Proposition 8.3 and we cannot apply the result of Proposition 8.3 in determining the changes in the drift of the Rec-ISQ-2 system due to the truncated stream. Letting $I^{r}$ denote the speed distribution in $\mathbb{P}(I^{r}=0)$, which now depends on the specific recycling stream $\{R_{t}\}$, which is difficult to characterize exactly.
@@
This is a characterization of the unused capacity in the Rec-ISQ-2 system, which is unaffected by the details of recycling stream. Therefore, our plan is to modify the affine-drift test function $h_{2}$ so that we get an unused-load term matching (9.1). Additionally, we use Lemma 7.6 instead of Lemma 7.5 because it supports recyclings. We start with the following test functions:
@@
In particular, we want to ensure that (9.2) is a constant not depending on $w$. By doing so, we will create a term matching the unused-load (9.1), allowing us to effectively bound the Rec-ISQ-2 system. We call this constant $C_{2}(x,\lambda_{x})$, though we do not yet know its exact value. The drift at speed $1/2$ is given by,
@@
We were able to apply the unused capacity formula (9.1) because our test function satisfied (9.2). This is key to the strength of our Rec-ISQ-2 bound Theorem 8.6.
@@ 10. Bounding Mean Relevant Work in the $M/G/k$
In this section, we extend the results in Section 8 for the ISQ-2 system to the general ISQ-$k$ system. The derivation of the test functions follows, the same ideas outlined in Section 9. We start by deriving our test functions using the DiffeDrift method. We first define the ISQ-$k$ constant-drift test function.
@@
$u_{q}(w)=e^{-\frac{kw\lambda}{q}}\int_{0}^{w}\dfrac{e^{\frac{k\lambda y}{q}}(k-q+k\lambda\mathbb{E}[u_{q+1}(S+y)])}{q}\,dy,$ (10.1)
@@
$v_{q}(w)=e^{-\frac{kw\lambda}{q}}\int_{0}^{w}\dfrac{e^{\frac{k\lambda y}{q}}(2ky-2qy+k\lambda\mathbb{E}[v_{q+1}(S+y)])}{q}\,dy,$ (10.2)
@@
Note that for $k=2$ this simplifies to the $h_{2}$ expression defined in Definition 8.2. Note also that the $v_{q}(w)$ formulas have a similar recursive structure as $u_{q}(w)$, but with an additional $y$ coefficient in the numerator of the integrand fraction. Equations 10.1 and 10.2 can be explicitly solved for any $k$, see Appendix C for the 3,4 and 5 server cases, and Section 11 for further discussion on $k>5$. In Lemmas B.1 and B.3 of Appendix B we prove the validity of these test functions.
@@
For an arbitrary job size $S$ such that $E[S^{3}]$ is finite and arrival rate $\lambda$, the expected total work in an ISQ-$k$ system is given by
@@
By Lemma B.2 and Lemma B.4 in Appendix B the drifts of the constant and affine test functions at speed 0 are given by $G\circ g_{k}(0,0)=\lambda\mathbb{E}[S+u_{1}(S)]$ and $G\circ h_{k}(0,0)=\lambda\mathbb{E}[S^{2}+v_{1}(S)]$ respectively. At all other speeds $i\geq 1/k$, the drift of the constant-drift test function is given by $G\circ g_{k}(w,i)=\lambda\mathbb{E}[S]-1$ and the drift of the affine-drift test function is given by $G\circ h_{k}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(\lambda\mathbb{E}[S]-1)$.
@@
Next, using Proposition 10.3 we give the exact mean relevant work in the Sep-ISQ-$k$ system. This constitutes the first novel relevant work lower bound of Theorem 4.2.
@@
For an arbitrary job size $S$ and arrival rate $\lambda$, the expected relevant work in the separate-ISQ-k system is given by,
$\mathbb{E}[W_{x}^{\text{Sep-ISQ-}k}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}+\dfrac{k\lambda(1-F_{S}(x))x^{2}}{2},$ (10.5)
where $v_{1}$ and $u_{1}$ are defined in Equations 10.1 and 10.2.
@@
The relevant work in the separate-ISQ-$k$ system is the sum of the total work in the truncated-ISQ-$k$ and the relevant work in the separate $M/G/\infty$ system.
@@
For the last part of this section, we derive the Rec-ISQ-$k$ lower bound. We start with the ISQ-$k$ modified affine-test function, denoted by $h_{k,x}$.
@@
$\ell_{q}(w)=e^{-\frac{kw\lambda_{x}}{q}}\int_{0}^{w}\dfrac{e^{\frac{k\lambda_{x}y}{q}}\left(k(q-k)C_{k}(x,\lambda_{x})+2(k-q)y+k\lambda_{x}\mathbb{E}[\ell_{q+1}(S_{x}+y)]\right)}{q}\,dy$ (10.6)
@@
where $v_{1}$ and $u_{1}$ are defined by Equations 10.1 and 10.2.
@@
For an arbitrary job size $S$ and arrival rate $\lambda$, a lower bound on expected relevant work $\mathbb{E}[W_{x}^{\text{Rec-ISQ-}k}]$ in the Rec-ISQ-$k$ system is given by,
$\mathbb{E}[W_{x}^{\text{Rec-ISQ-}k\text{-}L}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}\cdot\dfrac{1-\rho_{\bar{x}}}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})J_{x}}{2(1-\rho_{x})}.$ (10.8)
@@
We believe that $J_{x}=x^{2}$ for all thresholds $x$, numbers of servers $k$, and job size $S$. We proved in Theorem 8.6 that $J_{x}=x^{2}$ whenever $k=2$. However, formal proof of this assertion for $k\geq 3$ remains elusive, so we leave this as an open problem. We have empirically verified that this is true for the 3-server setting with exponential job sizes.
@@
For any job size $S$ and any number of servers $k\geq 2,J_{x}=x^{2}$.
10.1. Proof of Theorem 4.2
In this subsection, we prove Theorem 4.2.
Proof.
To prove Theorem 4.2 we combine results established in Section 5 and Section 10. We obtain Equation 4.6 by applying the WINE formula to Equations 5.2 and 5.3.
Equation 4.7 further incorporates $\mathbb{E}[W_{x}^{\text{Sep-ISQ-}k}]$. The fact that Sep-ISQ-$k$ lower bounds the relevant work of any $M/G/k$ system under arbitrary scheduling policy is proved in Theorem 6.2 and we derive its analytical expression given by Equation 4.9 in Theorem 10.4.
Finally, Equation 4.8 incorporates $\mathbb{E}[W_{x}^{\text{Rec-ISQ-}k}]$. The fact that there exists a Rec-ISQ-$k$ system lower bounding the relevant work of any $M/G/k$ system under arbitrary scheduling policy is proved in Theorem 6.3 and we derive a lower bound on $\mathbb{E}[W_{x}^{\text{Rec-ISQ-k}}]$ given by Equation 4.10 in Theorem 10.4.
11. Discussion on Generalization
In this section, we discuss how our theoretical framework naturally extends to the setting where arrivals have job sizes that are unknown but drawn from a known distribution, or where arrivals come with estimates. In addition, we address how to solve the recursive formulas Equations 10.1 and 10.2 for deriving ISQ-based lower bounds when $k$ is large.
11.1. Unknown size
Our framework can be extended to the setting where arrivals have job sizes that are unknown but drawn from a known distribution. The WINE formula, originally introduced by Scully2020, shows that the mean response time can be expressed as
$\mathbb{E}[T]=\frac{1}{\lambda}\int_{0}^{\infty}\frac{\mathbb{E}[W_{r}]}{r^{2}}\,dr,$ (11.1)
where $W_{r}$ denotes the total $r$-relevant Gittins work in the system. A job is said to be $r$-relevant if its Gittins rank is less than the cutoff $r$, and $W_{r}$ is computed by summing over all jobs, the service each requires before its Gittins rank exceeds $r$. Note that this $r$ is unrelated to the recycling state, and we are using standard notation.
The Gittins rank is a function of attained service $s$ and is defined as
$\text{Gittins-rank}(s)=\inf_{t>s}\frac{\mathbb{E}[\min\{S,t\}-s\mid S>s]}{\mathbb{P}(S\leq t\mid S>s)}.$ (11.2)
The Gittins rank was originally introduced by Gittins1979 to define the Gittins policy. The Gittins policy in scheduling selects the job with the smallest Gittins rank and is known to be optimal in the $M/G/1$ system. However, the WINE formula (11.1) is completely general, in particular, it holds for any scheduling policy, even those unrelated to the Gittins policy.
In the known size setting, the Gittins rank reduces to the remaining size, so $W_{r}$ is simply the sum of remaining work less than $r$, and Equation 11.1 reduces to Equation 5.1.
The MixEx bound carries over naturally into this setting. By computing the mean relevant work $\mathbb{E}[W_{r}]$ in the $M/G/\infty$ queue, together with the resource-pooled Gittins policy, we can plug these into the WINE framework to obtain the MixEx bound in the unknown job size setting, analogous to Equation 4.6. In addition, the Gittins-$k$ policy is the corresponding upper bound, which is heavy-traffic optimal in the unknown job size setting (Scully2020).
To illustrate the Gittins policy and Gittins rank, let us consider two specific cases where the job size distribution follows either a decreasing hazard rate (DHR) or an increasing hazard rate (IHR). In the DHR case, all jobs arrive to the system with their minimum Gittins rank, which is equal to the reciprocal of their hazard rate. Therefore, the Gittins policy reduces to the least-attained-service (LAS) policy and the Gittins work is easy to compute. In the IHR case, jobs arrive to the system with rank >0>0 and their rank gradually decreases over time until completion. In this case, the Gittins policy reduces to FCFS.
Next, we generalize the ISQ and ISQ-Recycling bounds in these two cases. In the DHR case, all jobs arrive to the system with their minimum rank and gradually increase over time until completion. Thus, there is no recycling in this system and our lower bound is given by applying the ISQ-$k$ formula to a truncated job size distribution.
In the IHR case, depending on the cutoff $r$ and the job size distribution, either all jobs enter the ISQ-$k$ system or all jobs recycle. Thus, ISQ only gives a nontrivial bound for ranks above the initial rank of the distribution and once again that lower bound is given by the ISQ-$k$ total work formula, Equation 10.3.
The Gittins rank function and the corresponding ISQ bounds are very simple in the IHR and DHR cases. For more complicated distributions, the Gittins rank function is non-trivial but related ISQ bounds can still be proven with additional effort.
11.2. Estimated size
Our framework can also incorporate the setting where arrivals come with estimates of job sizes. Just as in the case of unknown sizes, there are corresponding variants of the Gittins rank and corresponding WINE results in the setting where the joint distribution of size and estimate is known. For any such joint distribution, similar results to those outlined in Section 11.1 are possible.
When the estimates are of high quality, for jobs whose time in service is less than their estimate, the Gittins rank is approximately equal to the remaining size. Therefore, the analysis becomes more similar to the setting considered in this paper and the Gittins rank and relevant Gittins work are easier to compute, and thus similar ISQ bounds may be proven.
11.3. Computational considerations for large $k$.
In Appendix C, we explicitly derive the ISQ-$k$ relevant work lower bounds and response time lower bounds for $k=3,4,$ and $5$. We observe that the number of terms in these relevant work bounds roughly doubles as $k$ increases. Looking more carefully, the number of terms in $u_{i}$ and $v_{i}$, which are also given in Appendix C double per iteration as $i$ decreases. As a result, manually writing down and computing the numerical integrals for $k>5$ becomes difficult and a standard symbolic algebra package such as Mathematica cease to function effectively. However, one could implement a specialized computational program to compute these recursive functions, as defined in Equations 10.1 and 10.2.
In particular, the recursive equations involve only three types of terms: one constant term, one linear term in $w$, and many structured exponential terms of the form
$z_{c,r,m,[q]}(w)=c\cdot\frac{e^{-rw\lambda\prod_{i}\widetilde{S}(q_{i}\lambda)}}{\lambda^{m}},$ (11.3)
which can be characterized by a quadruple $(c,r,m,[q])$, where $c$ and $r$ are rationals, $m$ is a nonnegative integer, and $[q]$ denotes the list of rationals appearing in the product of transforms.
Given such a representation, each iteration can be efficiently performed in closed form. Specifically, each time we take integrals of the form
$u_{q}(w)=e^{-\tfrac{kw\lambda}{q}}\int_{0}^{w}\frac{e^{\tfrac{k\lambda y}{q}}\big(k\lambda\,\mathbb{E}[z_{c,r,m,[q]}(S+y)]\big)}{q}\,dy,$
as in Equations 10.1 and 10.2, the result produces two terms of exactly the same structure. Thus, it is feasible to build a computational tool to iteratively generate these terms and extend the analysis to $k\approx 20$, since the branching factor of the recursion is at most 2. We leave the implementation of such a computational program to future work.